18. Sparky Corporation uses the weighted-average method of process costing. The following information is available for February in its Molding Department:

Units:
Beginning Inventory: 31,000 units, 100% complete as to materials and 55% complete as to conversion.
Units started and completed: 122,000.
Units completed and transferred out: 153,000.
Ending Inventory: 33,000 units, 100% complete as to materials and 30% complete as to conversion.

Costs:
Costs in beginning Work in Process - Direct Materials: $49,000.
Costs in beginning Work in Process - Conversion: $54,850.
Costs incurred in February - Direct Materials: $293,000.
Costs incurred in February - Conversion: $605,150.

Calculate the equivalent units of materials.

19. Pitt Enterprises manufactures jeans. All materials are introduced at the beginning of the manufacturing process in the Cutting Department. Conversion costs are incurred uniformly throughout the manufacturing process. As the cutting of material is completed, the pieces are immediately transferred to the Sewing Department. Information for the Cutting Department for the month of May follows.

Work in Process, May 1 (56,500 units, 100% complete for direct materials, 40% complete with respect to conversion costs; includes $83,500 of direct material cost; $47,050 of conversion costs).

Units started in May 238,000
Units completed in May 213,000


Work in Process, May 31 (81,500 units, 100% complete for direct materials; 15% complete for conversion costs).

Costs incurred in May
Direct materials $ 423,640
Conversion costs $ 486,300


If Pitt Enterprises uses the FIFO method of process costing, compute the cost per equivalent unit for direct materials and conversion costs respectively for May.

20. At the beginning of the month, the Forming Department of Martin Manufacturing had 14,000 units in inventory, 40% complete as to materials, and 20% complete as to conversion. During the month the department started 68,000 units and transferred 70,000 units to the next manufacturing department. At the end of the month, the department had 12,000 units in inventory, 85% complete as to materials and 60% complete as to conversion. If Martin Manufacturing uses the FIFO method of process costing, compute the equivalent units for materials and conversion respectively for the Forming Department.
Answer #1

18) Equivalent unit of material (weighted average) = Unit transferred out+(Ending WIP*percent completion)

                                                                         = 153000+(33000*100%)

Equivalent unit of material (weighted average) = 186000

19) Equivalent unit of material (FIFO) = (56500*0%+156500*100%+81500*100%) = 238000

Equivalent unit of conversion (FIFO) = (56500*60%+156500*100%+81500*15%) = 202625

Cost per equivalent of material = 423640/238000 = 1.78

Cost per equivalent unit of conversion = 486300/202625 = 2.40

20) Equivalent unit of material = (14000*60%+56000*100%+12000*85%) = 74600

Equivalent unit of conversion = (14000*80%+56000*100%+12000*60%) = 74400